OpenStudy (anonymous):

A student measures two numbers x and y as x=10+-1 and y=20+-1 What is her best estimate for their product q=xy?

OpenStudy (fretje):

again the relative errors (faults) in a product need to be added: the relative errors are for x = r.f. = 1/10 = 0.1 for y = rf. = 1/20 = 0.05 added rf = 0.15 q = x.y = 200 with rf 0.15 so absolute error is 200*0.15 = +-30 so q = 200 +-30